OpenStudy (anonymous):

solutions for x^2+6x+1

OpenStudy (amoodarya):

\[x^2+6x+1 =(x+3)^2 -3^2 +1=(x+3)^2 -8\]

OpenStudy (tkhunny):

Well, you would need an equation for there to be any solutions. You should be able to find zeros with the quadratic formula or the lovely completing the square that has been demonstrated.